OpenStudy (anonymous):

please help. 5x-2y=-10 3x+6y=66 elimination by multiplacation or addition/subtraction.

OpenStudy (anonymous):

solve for y for either equations 5x-2y=-10 3x+6y=66 3x+6y =66 -3x -3x 6y = -3x + 66 divide both side by 6 y = -(1/2)x +11 plug y into 5x-2y=-10 5x-2(-(1/2)x+11)) = -10 5x + x - 22 = -10 + 22 +22 6x = 12 divide both side by 6 x = 2 Plug x into 3x+6y=66 3(2) + 6 y = 66 -6 -6 6y = 60 divide both side by y y=10 (2,10)

OpenStudy (anonymous):

opps I meant 6 the last part instead of y and check over the work in case I made a mistake somewhere

OpenStudy (anonymous):

thank you so much! i appreciate your help.

OpenStudy (anonymous):

your welcome!

OpenStudy (anonymous):

:)